Report of Activities on the RVIB N.B. Palmer
18 August 2001

A storm centered to the north and producing northerly winds of 50 kts at Palmer Station also keep us in cloud cover all of the yesterday, but because we were so much further south our winds were in the 20 kt range and out of the southwest. Weather this morning (20 August) is windy, colder, and not so cloudy. There was a nice pre-sunrise with shades of red highlighting the clouds where the sun will appear. At 0834, we are at -68° 44.410S; -74° 12.43W. Wind is out of 060 (east-northeast) at 22 to 27 kts, whipping up the recently fallen snow and lowering the ground visibility. The air temperature is down to -14.2°C and the barometer is 966.2 mlb.

On 18 August, only two stations (#73 and 74) were completed. This was in part due to a white-out condition that forced us to stay at station #73 for about 5 hours waiting for conditions to improve. It was also a day where extra time was needed to install an intermediate ice buoy at station #74. In addition, adding to the time demand was the fact that it has become more difficult and time consuming to attempt to make net tows. There were only two tasks scheduled for station #73, a CTD and a phytoplankton ring net. At station #74, in addition to the ice buoy deployment and ice collection, a CTD was done. An attempt to take a 10-m2 MOCNESS involved steaming the ship a distance of 4 nm into the wind and then returning along the trackline to the beginning point where the tow could be started. The hope was that the opening cut by the ship would remain open long enough to conduct the tow without having to back and ram. On this occasion, it was not to be. The ice floes were under a lot of pressure and the gap for much of the trackline was closed tight before the tow could be started and so it was cancelled. The steam from station #74 towards #75 was also difficult as the ice ridges occurred in increasing numbers as we approached the coastline of Alexander Island. There was a period during the day while steaming between stations #73 and 74 when a number of leads appeared and BIOMAPER-II was deployed for a good portion of the transit distance.

John Klinck reports on 18 August, the CTD group did two casts today at mid-shelf along the 180 line.

Station #73 (540 m) has a mixed layer to about 80 m, but there is some temperature and salinity structure over this layer. The pycnocline (100 to 250 m) has numerous temperature reversals although the layers are only a few meters thick. There is a slight O2 min at 350 m, but deep oxygen is almost uniform. Deep temperature and salinity are ever increasing below the pycnocline.

Station #74 (298 m) has a uniform mixed layer to 80 m. From 100 to 200 m, there are a series of steps in temperature and salinity with thicknesses of a few to 10 m or more. Steppy structure is evident to the bottom. No temperature maximum or O2 minimum occurs in the profile.

Jay Peterson reports that on work done at some stations on 17 August. At station #62, ADCP backscatter data indicated that there were concentrations of zooplankton at 70 m depth and a thick section from 120-190 m. SIMRAD data supported these findings and indicated an additional layer at ~300 m on the 38 KHz channel. Unfortunately, no OPC data was collected at station #62.

At station #63, an OPC tow was done in association with the MOC-1 to a depth of 240 m. Results from the OPC show that the greatest biovolume and abundance of zooplankton are found below 150 m, with slightly less in the upper water column. Below 150 m counts ranged from 750 - 1200 m-3 and above 150 m the range was 200 - 600 m-3. A plot of the counts-spectra showed a nearly linear decline in numbers with an increase in size category for all nets. A slight deviation in this pattern occurred at the size category of large copepods and small furcilia. Visual observation of the contents of each net revealed a domination of Thysanoessa sp. juveniles and late stage Calanoid copepods. ADCP backscatter data collected during the tow showed scattering layers at 40 m and 160 m.

Frank Stewart has provided an update on the ice collecting work on board the Palmer. This report concerns work done at broad-scale station #74. Sea ice biology core #27 was taken concurrent with the deployment of ARGOS intermediate buoy #7950 at grid station #74 (68° 39.39S 72° 56.51W) on 18 August (JD 230). The core site was located within 15 m of the buoy on a flat section of a vast floe of first year sea ice. The floe was weighted heavily by >30 cm of fresh snow that depressed the ice surface below freeboard in some areas and produced a well-defined slush layer at the snow-ice interface. Ice at the core hole itself was topped with 38cm of snow, had a negative freeboard of 1cm (water line 1cm above ice surface), and was 100 cm thick (as measured with ice thickness tape). Following extraction with a hand-held Kovacs core barrel, the core was divided and sub-sampled into four sections based on natural breakage points in the ice: sections 0-14 cm depth, 15-23 cm, 24-85 cm, and 86-100 cm. The top two sections and the bottom section were of similar texture and coloration: relatively soft, i.e., high-liquid water content, and gray. The long middle section (24-85 cm) was of well-consolidated, hard gray ice with intermittent bands of whiter ice. A 12 cm-section of tightly packed snow/slush that remained in the core barrel atop the ice layers was also sampled. The brine/seawater that infiltrated the core hole after the core was removed was sampled using a Nalgene bottle attached to a meter stick. Samples were brought on board and ice core sections, including the packed snow layer, were diluted with 0.2 mm-filtered seawater at a ratio of 2:1 sea water:core meltwater. The brine sample was not diluted. All samples were left to melt in the dark at 4-7°C. Melted sub-samples will be processed to determine rates of bacterial production (as measured by uptake of tritiated thymidine over time) at 0°C and 4°C, bacterial and algal biomass, and dissolved organic carbon concentrations. Fractions of the sub-samples will be consolidated, filtered, and preserved for electrophoretic and PCR-based analysis of the taxonomic composition of the ice bacterial community. In at least one of the sub-samples, production by the autotrophic component of the ice community will be measured by the uptake of carbon-14 over a range of light levels (photosynthesis vs. irradiance curve courtesy of W. Kozlowski). Core 27 is similar in terms of structure and length to Core 17, a highly productive and high biomass core taken from a field of consolidated first-year cake and small floes off the shelf break at station #47. The thickness and internal variability of both cores are indicative of ice that has formed over several months, and potentially over the course of several flood-freeze cycles, and therefore may be harboring the remnants of a fall "bloom" in the ice microbial community.

Ari Friedlaender reports that on 18 August observations began as the Palmer steamed to station #74 at 0930. Conditions were very poor with less than 0.5 nm visibility due to snow fog. Ice coverage was 10/10 with some narrow, yet extensive leads that were covered in grease ice or new nilas for the most part. We arrived at station at 1500 (-68° 39.326S, -72° 57.824W). One crabeater seal was seen in a lead, and no whales were observed.

Ana Sirovic reports that she deployed 1 sonobuoy, 15 nm from station 74 into a lead from the port side of the ship. As the ice in the area was not heavy, reception was good for a long time: 2h 29 min. During this time, some seal calls were heard (downsweeps around 300 Hz). That was the only deployment for the day due the ice buoy installation and coincidentally, the football game on the ice.

Chris Ribic and Erik Chapman report that on 18 August (JD-230), they surveyed surveyed for 4 hours and 30 minutes between stations #73 and 74. Visibility was poor and the 300 m transect was only visible for 2 hours during the transit. The ship traveled through leads to pick through the heavy ice. Since birds are mainly seen in association with leads, the survey probably overestimates the number of birds in the area. Even so, few birds were seen including a single Adélie Penguin and Antarctic Petrel in addition to 12 Snow Petrels, most of which were found in association with a single large lead. These numbers were surprisingly low, considering how much time was spent in leads today. A summary of their observations is the following:

Primary Ice: Mostly vast floes. Some new gray ice in leads, a lots of open water in leads.
Common Name Number
Antarctic Petrel 1
Adélie Penguin 1
Snow Petrel 12

BIOMAPER-II/MOCNESS report (P. Wiebe, C. Ashjian, and S. Gallager):
BIOMAPER-II was deployed on 18 August about 1030 after leaving station #73. The Palmer had entered an area with a substantial number of leads and was steaming towards station #74 by moving through one lead and then steering over to another. There was one major hangup that Christian MacDonald caught, in which the ship needed to back down and blow the ice out from around the wire, while the fish sank down to 277 m. And there was another short snag. But all in all, it was good towyoing country. We pulled BIOMAPER-II out of the water just a short distance from station #74. Somehow, we managed to keep it in the water for almost the entire 20-mile section, but towards the end, the going got pretty iffy and there were a number of stops to back and ram.

There was a thin backscattering layer at about 140 m depth and in the zone above to the surface, there was almost no scattering on the 120 or 200 kHz echograms. Much more intense scattering occurred in a deeper layer that started around 250 m. When the wire hung up on an ice chunk and the ship stopped, the towed body dipped down into the higher backscattering layer and the VPR captured images of the animals there, which were dominated by copepods.

No 1-m2 MOCNESS tow was done on 18 August.

Cheers, Peter